OpenStudy (anonymous):

Talia took 15 test and got an average of 73. After dropping her lowest score, her average went up to 74.5. What was her dropped score?

OpenStudy (ganpat):

avg for 15 = 73.. so total score = 15 * 73 = 1095 avg after dropping lowest score = 74.5 so total of 14 scores = 74.5 * 14 = 1043 lowest score = 1095 - 1043 = 52

OpenStudy (anonymous):

Thanks VERY MUCH big thanks to you